What is the equation of an ellipse with center (-4,2) and a horizontal major axis?
Answer by venugopalramana(3286)

What is the equation of an ellipse with center (-4,2)=(H,K) and a horizontal major axis?
IT DEPENDS ON LENGTH OF MAJOR AND MINOR AXES...IF THEY ARE 2A AND 2B THEN EQN.IS
(X-H)^2/A^2+(Y-K)^2/B^2=1
(X+4)^2/A^2+(Y-2)^2/B^2=1